# DST-Adapter: Disentangled-and-Deformable Spatio-Temporal Adapter for Few-shot Action Recognition

### v1

@@ Abstract
Adapting large pre-trained image models to few-shot action recognition has proven to be an effective and efficient strategy for learning robust feature extractors, which is essential for few-shot learning. Typical fine-tuning based adaptation paradigm is prone to overfitting in the few-shot learning scenarios and offers little modeling flexibility for learning temporal features in video data. In this work we present the Disentangled-and-Deformable Spatio-Temporal Adapter (DST-Adapter), a novel adapter tuning framework for few-shot action recognition, which is designed in a dual-pathway architecture to encode spatial and temporal features in a disentangled manner. Furthermore, we devise the Deformable Spatio-Temporal Attention module as the core component of DST-Adapter, which can be tailored to model both spatial and temporal features in corresponding pathways, allowing our DST-Adapter to encode features in a global view in 3D spatio-temporal space while maintaining a lightweight design. Extensive experiments with instantiations of our method on both pre-trained ResNet and ViT demonstrate the superiority of our method over state-of-the-art methods for few-shot action recognition. Our method is particularly well-suited to challenging scenarios where temporal dynamics are critical for action recognition.
1 Introduction
@@
As a prominent parameter-efficient fine-tuning technique [14, 56, 17, 12], adapter tuning [14, 4, 39] embeds learnable adapters with a negligible amount of parameters into pre-trained models flexibly while keeping the model frozen, conducting task adaptation efficiently and effectively. ST-Adapter [25] makes the first attempt to apply adapter tuning to image-to-video transfer learning. As shown in Figure 1, it follows the classical bottleneck structure for adapter design [14] and inserts a 3D convolution layer in the bottleneck for feature transformation. Such a simple design can learn spatio-temporal features while maintaining a lightweight structure. Nevertheless, we investigate two potential limitations of ST-Adapter. First, it learns the spatial and temporal features jointly by 3D convolution, whereas it has been shown that such a joint learning paradigm of spatio-temporal features for video data is inferior to learning the spatial and temporal features in a disentangled way, such as SlowFast [8] or two-stream design [29, 38]. Second, the convolutional operation, especially with shallow layers for lightweight design, has a limited local receptive field in both the spatial and temporal domains, which also limits the performance of ST-Adapter.
@@
In this work we design the Disentangled-and-Deformable Spatio-Temporal Adapter (DST-Adapter) to tackle the limitations of ST-Adapter. Specifically, we design our DST-Adapter as a dual-pathway architecture, as illustrated in Figure 1, in which the spatial pathway is responsible for capturing the spatial appearance features while the temporal pathway focuses on learning the temporal dynamics. Thus, our model can encode the spatio-temporal features in a disentangled manner. Moreover, we devise the Deformable Spatio-Temporal Attention (DSTA) module as an essential component to model both the spatial and temporal pathways. It adapts the deformable attention [46] from 2D image space to 3D spatio-temporal space, which first samples a group of reference points in the spatio-temporal space and then learns the offset for each point to shift them to more informative regions in the 3D space. These shifted points are used as key and value pairs for DSTA to perform sparse self-attention, which allows our DST-Adapter to perform feature adaptation in a global view while maintaining a lightweight design.
@@
A novel design of our DSTA is that we make the sampling density along the spatial and temporal domains configurable, which allows for tailoring specialized versions of DSTA to model the spatial and temporal pathways separately. In particular, we tailor DSTA-T with denser sampling along the temporal domain than the spatial domain for the temporal pathway since it focuses on capturing temporal features. In contrast, the tailored DSTA-S for the spatial pathway samples denser reference points along the spatial domain. To conclude, we make the following contributions:
@@
We propose DST-Adapter, a novel adapter tuning framework for few-shot action recognition, which is designed in a dual-pathway architecture to encode the spatial and temporal features in a disentangled manner.
@@
We devise the Deformable Spatio-Temporal Attention (DSTA), which can be tailored to model both the spatial and temporal pathways, allowing our DST-Adapter to encode features in a global view while maintaining a lightweight design.
@@
Extensive experiments on five standard benchmarks with instantiations of our method on pre-trained ResNet [13] and ViT [27], demonstrate the superiority of our method over other state-of-the-art methods, particularly in challenging scenarios like the SSv2 benchmark where the temporal dynamics are critical for action recognition. Besides, comprehensive ablation study is conducted to investigate our DST-Adapter thoroughly.
@@ 2 Related Work
Few-shot action recognition. Most existing few-shot action recognition methods adopt the metric-based paradigm to classify videos, and primarily focus on two directions to deal with this task. The first one is to investigate the spatio-temporal modeling [57, 58, 52, 44, 54, 34, 42, 48, 45]. For example, MTFAN [44] proposes a motion encoder to learn global motion patterns and injects them into each video representations using a motion modulator. STRM [34] enriches local patch features and global frame features for joint spatio-temporal modeling. MoLo [42] designs a motion autodecoder to explicitly extract motion dynamics in a unified network. Our method also aims to model spatio-temporal features effectively and efficiently based on adapter-tuning technique. Another direction focuses on designing effective metric learning strategy [2, 1, 26, 40, 44, 15, 24, 55]. OTAM [2] utilizes the DTW [23] algorithm to calculate video distances with strict temporal alignment. TRX [26] exhaustively enumerates all sub-sequences of support and query videos and matches them using an attention mechanism. HyRSM [40] applies a novel bidirectional mean Hausdorff metric (referred to as Bi-MHM) to alleviate the strictly ordered constraints. We evaluate our DST-Adapter using the above three classical matching strategies to demonstrate its effectiveness and robustness.
Adapter tuning. As a classical parameter-efficient fine-tuning method, adapter tuning method is first proposed in [14], and quickly draws attention in many other research areas [4, 39, 12, 33]. A common practice is to build up a lightweight module (named Adapter) which only consisting of negligible learnable parameters, and selectively plug it into a pre-trained model. During training, only the parameters of inserted adapters are updated while the original model remains frozen, leading to efficient task adaptation. Recently, some works apply this method to adapt image models for video recognition. AIM [49] inserts adapters into different parts of each ViT [7] stage for spatial, temporal, and joint adaptation, separately. ST-Adapter [25] employs depth-wise 3D convolution to construct adapter for feature adaptation, which endows it with spatio-temporal modeling capability. Our DST-Adapter follows the basic framework of ST-Adapter, and meanwhile optimize the essential technical designs for few-shot action recognition.
@@ 3.1 Overview
Problem Formulation. The aim of few-shot action recognition is to learn an action recognizer from a set of base classes, which is able to recognize novel classes using only a few labeled samples. The task typically follows episodic paradigm [37, 2, 40, 42], which is consistent with general application scenarios for few-shot learning. An episode includes a support set consisting of classes and labeled samples for each class (referred to as the -way -shot task), as well as a query set that contains unlabeled samples to be classified. In each episode, we aim to classify every query into one of the classes with the guidance of the support set. Such episodic task setting is consistently adopted during all the training, validation, and test stages.
@@
Adapter Tuning Framework. We design a plug-and-play and lightweight adapter, dubbed DST-Adapter, which can be integrated into most of existing deep feature learning framework of vision models. Thus we can leverage the powerful feature encoding capability of pre-trained large vision models by adapting them into the task of few-shot action recognition efficiently while incurring only a small amount of parameter-tuning overhead.
@@
Figure 2 (a) illustrates the overall adapter tuning framework of our method. Given a pre-trained feature extractor from a large model, our designed lightweight DST-Adapter can be selectively plugged into the middle layers of the feature extractor to perform feature transformation for task adaptation while maintaining the shape of feature maps consistent before and after each adaptation. Only the inserted adapters are tuned during training while the pre-trained backbone keeps frozen, leading to efficient task adaptation. The learned features by such adapter tuning framework are further used for few-shot action recognition based on the metric-based strategy [30, 2, 42].
@@
Our DST-Adapter is designed independently from the backbone structure of the feature extractor, thus it can be readily integrated into most of classical pre-trained deep vision models. In particular, we instantiate the backbone of feature extractor with two dominant deep visual learning frameworks, namely ResNet [13] and ViT [7], respectively.
3.2 DST-Adapter
@@
Dual-pathway Adapter Architecture for Spatio-Temporal Disentanglement. Successful action recognition from video data entails effective feature learning of both spatial semantic and temporal dynamic features. Most large vision models are typically pre-trained on image data, thus our DST-Adapter should be capable of capturing both the spatial features and temporal features. To this end, we design the DST-Adapter as a dual-pathway architecture shown in Figure 2 (c), in which the spatial pathway is responsible for capturing the spatial semantics while the temporal pathway focuses on learning the temporal dynamics. As a result, our model is able to encode the spatio-temporal features for video data in a disentangled manner.
@@
As a common practice of typical Adapters [14, 25], our DST-Adapter adopts the bottleneck architecture for reducing the computational complexity. It first downsamples the feature map into a low-dimensional feature space, then the downsampled features are fed into the spatial and temporal pathways concurrently to perform disentangled feature adaptation. Finally, both the adapted spatial and temporal features are fused by simple element-wise addition and upsampled back to the initial size. Formally, given the feature maps obtained from the -th stage of the pre-trained backbone, containing channels of feature maps with spatial size for each of frames, the feature adaptation by DST-Adapter can be formulated as:
@@
Herein, and are the transformation matrices of two linear layers for downsampling and upsampling, respectively. and denote the disentangled feature adaptation by the spatial and temporal pathways respectively. Below we discuss two feasible modeling mechanisms for and , which leverages 3D convolution and our proposed Deformable Spatio-Temporal Attention, respectively.
@@
While it has been validated that leveraging 3D convolutions with different kernel shapes can learn the spatial and temporal features separately [8, 36], An important limitation of using 3D convolutions to construct DST-Adapter, which is also suffered by ST-Adapter, is that the convolutional operation has limited local receptive field either in the spatial or the temporal domain. Constructing deep convolutional network by stacking convolutional layers can enlarge the receptive field whilst increasing the model size and incurring heavy computational burden, which is inconsistent with the lightweight design principle for adapters.
@@
Modeling with Deformable Spatio-Temporal Attention. To address the limitation of the modeling mechanism based on 3D convolution (DST-Adapter), we devise Deformable Spatio-Temporal Attention (DSTA) which performs feature adaptation in a global view by sparse self-attention while maintaining high computational efficiency. It adapts the deformable attention [46] from 2D image space to 3D spatio-temporal space. As shown in Figure 2 (d), similar to the deformable attention, DSTA first samples a group of reference points in the spatio-temporal space and then learns the offset for each reference point to shift them to the more informative regions. As a result, DSTA can learn a set of informative tokens which are further used as key and value pairs shared for all queries for sparse self-attention.
@@
We conduct several adaptations from the deformable attention in 2D space to our DSTA in spatio-temporal 3D space. First, we employ Dynamic Position Embedding (DPE) [21] implemented as a 3D depth-wise convolution layer to learn the semantic-conditioned spatio-temporal position information for each token and incorporate it into the semantic features by simple element-wise addition. Thus, the feature map fused with the position information at the -th stage is obtained by:
@@
Our DSTA learns the offsets in 3D space for the reference points by a 3D convolutional network consisting of a 3D depth-wise convolution layer and a 3D convolution layer as well as a GELU in between. After shifting the reference points according to the learned offsets, we derive the features for each shifted reference point by performing trilinear interpolation among neighboring tokens within a 3D volume space around the point. For instance, the features for the shifted point located at is calculated via the trilinear interpolation by:
@@
where defines the interpolating volume space and indexes all token in the whole spatio-temporal 3D space. All shifted reference points are used as the keys and values shared for sparse self-attention. For instance, a token in located at serves as a query and the output of DSTA is calculated by:
@@
where is the stacked feature matrix of all shifted points and are projection matrices for the query, key, value and output respectively.
@@
The learned offsets for reference points is restricted in a limited range for two reasons. First, it can prevent the potential training collapse that all reference points are shifted to the same point. Second, such restriction can ensure that each reference point can be shifted to a unique location within a local region. As a result, more sampling of reference points typically lead to more powerful and fine-grained representation by the shifted points. The deformable attention samples reference points uniformly in the 2D feature space with the same sampling density along two spatial dimensions. In contrast, we make the sampling density in the spatial and temporal domains configurable. Consequently, our DST-Adapter can leverage DSTA to model both in the spatial pathway for spatial feature adaptation and in the temporal pathway for temporal feature adaptation. Specifically, we define the ‘Sampling kernel’ for DSTA as:
Definition 1.
@@
Sampling kernel: formulated as the sampling densities of reference points along the dimensions of time, height and width, respectively, denoted as , s.t. . Herein, is the total number of sampled reference points.
@@
Intuitively, a DSTA module whose sampling kernel has large and small are more capable of learning the temporal features than learning the spatial features. On the other hand, sampling more reference points in the spatial domain than the temporal domain () generally makes DSTA focus on learning the spatial features. Thus, we can tailor DSTA by configuring the sampling kernel to model and correspondingly. In particular, we tailor two versions of DSTA:
@@
DSTA-S for modeling in the spatial pathway, whose sampling kernel has relatively large and small ().
@@
DSTA-T for modeling in the temporal pathway, which samples denser reference points in the temporal domain than in the spatial domain ().

### v2

# D2ST-Adapter: Disentangled-and-Deformable Spatio-Temporal Adapter for Few-shot Action Recognition

@@ Abstract
Adapting large pre-trained image models to few-shot action recognition has proven to be an effective and efficient strategy for learning robust feature extractors, which is essential for few-shot learning. Typical fine-tuning based adaptation paradigm is prone to overfitting in the few-shot learning scenarios and offers little modeling flexibility for learning temporal features in video data. In this work we present the Disentangled-and-Deformable Spatio-Temporal Adapter (D2ST-Adapter), which is a novel adapter tuning framework well-suited for few-shot action recognition due to lightweight design and low parameter-learning overhead. It is designed in a dual-pathway architecture to encode spatial and temporal features in a disentangled manner. In particular, we devise the anisotropic Deformable Spatio-Temporal Attention module as the core component of D2ST-Adapter, which can be tailored with anisotropic sampling densities along spatial and temporal domains to learn spatial and temporal features specifically in corresponding pathways, allowing our D2ST-Adapter to encode features in a global view in 3D spatio-temporal space while maintaining a lightweight design. Extensive experiments with instantiations of our method on both pre-trained ResNet and ViT demonstrate the superiority of our method over state-of-the-art methods for few-shot action recognition. Our method is particularly well-suited to challenging scenarios where temporal dynamics are critical for action recognition.
Keywords:
Action recognition Few-shot learning Adapter tuning1 Introduction
@@
As a prominent parameter-efficient fine-tuning technique [14, 58, 17, 12], adapter tuning [14, 4, 41] embeds learnable adapters with a negligible amount of parameters into pre-trained models flexibly while keeping the model frozen, conducting model adaptation efficiently and effectively. Compared to the fine-tuning adaptation paradigm, adapter tuning is particularly better-suited to few-shot learning scenarios. A straightforward way to apply adapter tuning to image-to-video transfer learning is to employ the vanilla adapter for image data directly to adapt the pre-learned spatial features and temporal features separately, which is exactly adopted by AIM [51] and DUALPATH [26]. As shown in Figure 1, both AIM and DUALPATH first pre-learn the spatial and temporal features separately by duplicating the multi-head self-attention modules either in cascaded (AIM) or parallel (DUALPATH) manners, and insert vanilla adapter in each duplicated module for specific feature adaptation. However, such methods inevitably incur heavyweight model architecture and massive computation overhead. ST-Adapter [25] first attempts to design the specialized adapter for video data by inserting a 3D convolution layer into the vanilla adapter for learning spatio-temporal features while maintaining a lightweight structure. Nevertheless, we investigate two potential limitations of ST-Adapter. First, it learns the spatial and temporal features jointly by 3D convolution, whereas it has been shown that in the low-data scenarios, such a joint learning paradigm of spatio-temporal features for video data is inferior to learning the spatial and temporal features in a disentangled way, such as SlowFast [8] or two-stream design [31, 40]. Second, the convolutional operation, especially with shallow layers for lightweight design, has a limited local receptive field in both the spatial and temporal domains, which also limits the performance of ST-Adapter.
@@
In this work we design the Disentangled-and-Deformable Spatio-Temporal Adapter (D2ST-Adapter) to tackle the aforementioned limitations. Specifically, we design our D2ST-Adapter as a dual-pathway architecture, as illustrated in Figure 1, in which the spatial pathway is responsible for capturing the spatial appearance features while the temporal pathway focuses on learning the temporal dynamics. Thus, our model can encode the spatio-temporal features in a disentangled manner. Moreover, we devise the anisotropic Deformable Spatio-Temporal Attention (aDSTA) module as an essential component to model both the spatial and temporal pathways. It adapts the deformable attention [48] from 2D image space to 3D spatio-temporal space, which first samples a group of reference points in the spatio-temporal space and then learns the offset for each point to shift them to more informative regions in the 3D space. These shifted points are used as key and value pairs for aDSTA to perform sparse self-attention, which allows our D2ST-Adapter to perform feature adaptation in a global view while maintaining a lightweight design.
@@
A novel design of our aDSTA is that we make the sampling density anisotropic along the spatial and temporal domains by specific configuration, which allows for tailoring specialized versions of aDSTA to model the spatial and temporal pathways separately. In particular, we tailor aDSTA-T with denser sampling along the temporal domain than the spatial domain for the temporal pathway since it focuses on capturing temporal features. In contrast, the tailored aDSTA-S for the spatial pathway samples denser reference points along the spatial domain. To conclude, we make the following contributions:
@@
We propose D2ST-Adapter, a novel adapter tuning method for few-shot action recognition, which is designed in a dual-pathway architecture to adaptively encode the spatial and temporal features in a disentangled manner. It can adapt large vision models pre-trained on image data to video data efficiently and effectively in few-shot learning scenarios due to the lightweight design.
@@
We devise the anisotropic Deformable Spatio-Temporal Attention (aDSTA), which can be tailored with anisotropic sampling densities along spatial and temporal domains to model the spatial and temporal pathways separately, allowing our D2ST-Adapter to encode features in a global view while maintaining a lightweight design.
@@
Extensive experiments on five benchmarks with instantiations of our method on pre-trained ResNet [13] and ViT [29], demonstrate the superiority of our method over other methods, particularly in challenging scenarios like SSv2 benchmark where the temporal dynamics are critical for action recognition.
@@ 2 Related Work
Few-shot action recognition. Most existing few-shot action recognition methods adopt the metric-based paradigm to classify videos, and primarily focus on two directions to deal with this task. The first one is to investigate the spatio-temporal modeling [59, 60, 54, 27, 46, 56, 36, 43, 50, 47]. For example, MTFAN [46] proposes a motion encoder to learn global motion patterns and injects them into each video representations using a motion modulator. STRM [36] enriches local patch features and global frame features for joint spatio-temporal modeling. MoLo [43] designs a motion autodecoder to explicitly extract motion dynamics in a unified network. Our method also aims to model spatio-temporal features effectively and efficiently based on adapter-tuning technique. Another direction focuses on designing effective metric learning strategy [2, 1, 28, 44, 46, 15, 24, 57]. OTAM [2] utilizes the DTW [23] algorithm to calculate video distances with strict temporal alignment. TRX [28] exhaustively enumerates all sub-sequences of support and query videos and matches them using an attention mechanism. HyRSM [44] applies a novel bidirectional mean Hausdorff metric (referred to as Bi-MHM) to alleviate the strictly ordered constraints. We evaluate our D2ST-Adapter using the above three classical matching strategies to demonstrate its effectiveness and robustness.
Adapter tuning. As a classical parameter-efficient fine-tuning method, adapter tuning is first proposed in [14], and quickly draws attention in many other research areas [4, 41, 12, 35]. A common practice is to build up a lightweight module (named Adapter) which only consisting of negligible learnable parameters, and selectively plug it into a pre-trained model. During training, only the parameters of inserted adapters are updated while the original model remains frozen, leading to efficient task adaptation. Recently, some works apply this method to adapt image models for video recognition. AIM [51] duplicates multi-head self-attention modules and plugs adapters after them to separately learn the spatial and temporal features in a cascaded manner. Similarly, DUALPATH [26] explicitly builds the two-stream architecture upon ViT and utilizes adapters to parallelly learn the spatial and temporal features. ST-Adapter [25] employs depth-wise 3D convolution to construct adapter for feature adaptation, which endows it with spatio-temporal modeling capability. Our D2ST-Adapter follows the basic framework of ST-Adapter, and meanwhile optimize the essential technical designs for few-shot action recognition.
@@ 3.1 Overview
Problem Formulation of Few-shot Action Recognition. The aim of few-shot action recognition is to learn an action recognizer from a set of base classes, which is able to recognize novel classes using only a few labeled samples. The task typically follows episodic paradigm [39, 2, 44, 43], which is consistent with general application scenarios for few-shot learning. An episode includes a support set consisting of classes and labeled samples for each class (referred to as the -way -shot task), as well as a query set that contains unlabeled samples to be classified. In each episode, we aim to classify every query into one of the classes with the guidance of the support set. Such episodic task setting is consistently adopted during all the training, validation, and test stages.
@@
Adapter Tuning Framework. We design a plug-and-play and lightweight adapter, dubbed D2ST-Adapter, which can be integrated into most existing large vision models. Thus we can leverage the powerful feature encoding capability of pre-trained large vision models by efficient image-to-video model adaptation while incurring only a small amount of parameter-tuning overhead. It is particularly well-suited to few-shot learning scenarios.
@@
Figure 2 (a) illustrates the overall adapter tuning framework of our method. Given a pre-trained feature extractor from a large model, our designed lightweight D2ST-Adapter can be selectively plugged into the middle layers of the feature extractor to perform feature transformation for task adaptation while maintaining the shape of feature maps consistent before and after each adaptation. Only the inserted adapters are tuned during training while the pre-trained backbone keeps frozen, leading to efficient task adaptation. The learned features by such adapter tuning framework are further used for few-shot action recognition based on the metric-based strategy [32, 2, 43].
@@
Our D2ST-Adapter is designed independently from the backbone structure of the feature extractor, thus it can be readily integrated into most of classical pre-trained deep vision models. In particular, we instantiate the backbone of feature extractor with two dominant deep visual learning frameworks, namely ResNet [13] and ViT [7], respectively.
3.2 D2ST-Adapter
@@
Dual-pathway Adapter Architecture for Spatio-Temporal Disentanglement. Successful action recognition from video data entails effective feature learning of both spatial semantic and temporal dynamic features. Most large vision models are typically pre-trained on image data, thus our D2ST-Adapter should be capable of capturing both the spatial features and temporal features. To this end, we design the D2ST-Adapter as a dual-pathway architecture shown in Figure 2 (c), in which the spatial pathway is responsible for capturing the spatial semantics while the temporal pathway focuses on learning the temporal dynamics. As a result, our model is able to encode the spatio-temporal features for video data in a disentangled manner.
@@
As a common practice of typical Adapters [14, 25], our D2ST-Adapter adopts the bottleneck architecture for reducing the computational complexity. It first downsamples the feature map into a low-dimensional feature space, then the downsampled features are fed into the spatial and temporal pathways concurrently to perform disentangled feature adaptation. Finally, both the adapted spatial and temporal features are fused by simple element-wise addition and upsampled back to the initial size. Formally, given the feature maps obtained from the -th stage of the pre-trained backbone, containing channels of feature maps with spatial size for each of frames, the feature adaptation by D2ST-Adapter can be formulated as:
@@
Herein, and are the transformation matrices of two linear layers for downsampling and upsampling, respectively. and denote the disentangled feature adaptation by the spatial and temporal pathways respectively. Below we discuss two feasible modeling mechanisms for and , which leverages 3D convolution and our proposed Anisotropic Deformable Spatio-Temporal Attention, respectively.
@@
While it has been validated that leveraging 3D convolutions with different kernel shapes can learn the spatial and temporal features separately [8, 38], An important limitation of using 3D convolutions to construct D2ST-Adapter, which is also suffered by ST-Adapter, is that the convolutional operation has limited local receptive field either in the spatial or the temporal domain. Constructing deep convolutional network by stacking convolutional layers can enlarge the receptive field whilst increasing the model size and incurring heavy computational burden, which is inconsistent with the lightweight design principle for adapters.
@@
Modeling with Anisotropic Deformable Spatio-Temporal Attention. To address the limitation of the modeling mechanism based on 3D convolution (DST-Adapter), we devise anisotropic Deformable Spatio-Temporal Attention (aDSTA) which performs feature adaptation in a global view by sparse self-attention while maintaining high computational efficiency. It adapts the deformable attention [48] from 2D image space to 3D spatio-temporal space. As shown in Figure 2 (d), aDSTA first samples a group of reference points with anisotropic sampling density in the spatio-temporal space and then learns the offset for each reference point to shift them to the more informative regions. As a result, aDSTA can learn a set of informative tokens which are further used as key and value pairs shared for all queries for sparse self-attention.
@@
We conduct several adaptations from the deformable attention in 2D space to our aDSTA in spatio-temporal 3D space. First, we employ Dynamic Position Embedding (DPE) [21] implemented as a 3D depth-wise convolution layer to learn the semantic-conditioned spatio-temporal position information for each token and incorporate it into the semantic features by simple element-wise addition. Thus, the feature map fused with the position information at the -th stage is obtained by:
@@
Our aDSTA learns the offsets in 3D space for the reference points by a 3D convolutional network consisting of a 3D depth-wise convolution layer and a 3D convolution layer as well as a GELU in between. After shifting the reference points according to the learned offsets, we derive the features for each shifted reference point by performing trilinear interpolation among neighboring tokens within a 3D volume space around the point. For instance, the features for the shifted point located at is calculated via the trilinear interpolation by:
@@
where defines an interpolating cubic volume space centered at and indexes all tokens in the whole spatio-temporal 3D space. All shifted reference points are used as the keys and values shared for sparse self-attention. For instance, a token in located at serves as a query and the output of aDSTA is calculated by:
@@
where is feature matrix of shifted points and are projection matrices for the query, key, value and output respectively.
@@
Anisotropic Sampling Density. The learned offsets for reference points are restricted in a limited range for two reasons. First, it can prevent the potential training collapse that all reference points are shifted to the same point. Second, such restriction can ensure that each reference point can be shifted to a unique location within a local region. As a result, more sampling of reference points typically leads to more powerful and fine-grained representations by the shifted points. The deformable attention samples reference points uniformly in the 2D feature space with the same sampling density along two spatial dimensions. In contrast, we configure anisotropic sampling density in the spatial and temporal domains. Consequently, our D2ST-Adapter can leverage aDSTA to model both spatial feature adaptation in the spatial pathway and temporal feature adaptation in the temporal pathway. Specifically, we define the ‘Sampling kernel’ for aDSTA as:
Definition 1
@@
Sampling kernel: formulated as the sampling densities of reference points along the dimensions of time, height and width, respectively, denoted as , s.t. . is the total number of sampled reference points.
@@
Intuitively, an aDSTA module whose sampling kernel has large and small are more capable of learning the temporal features than learning the spatial features. On the other hand, sampling more reference points in the spatial domain than the temporal domain () generally makes aDSTA focus on learning the spatial features. Thus, we can tailor aDSTA by configuring the sampling kernel to model and correspondingly. In particular, we tailor two versions of aDSTA:
@@
aDSTA-S for modeling in the spatial pathway, whose sampling kernel has relatively large and small ().
@@
aDSTA-T for modeling in the temporal pathway, which samples denser reference points in temporal domain than in spatial domain ().